# MagicDrive-V2: High-Resolution Long Video Generation for Autonomous Driving with Adaptive Control

### v3

@@ Abstract
The rapid advancement of diffusion models has greatly improved video synthesis, especially in controllable video generation, which is vital for applications like autonomous driving. Although DiT with 3D VAE has become a standard framework for video generation, it introduces challenges in controllable driving video generation, especially for geometry control, rendering existing control methods ineffective. To address these issues, we propose MagicDrive-V2, a novel approach that integrates the MVDiT block and spatial-temporal conditional encoding to enable multi-view video generation and precise geometric control. Additionally, we introduce an efficient method for obtaining contextual descriptions for videos to support diverse textual control, along with a progressive training strategy using mixed video data to enhance training efficiency and generalizability. Consequently, MagicDrive-V2 enables multi-view driving video synthesis with $\times3$ resolution and 4×4$\times$4 × frame count (compared to current SOTA), rich contextual control, and geometric controls. Extensive experiments demonstrate MagicDrive-V2’s ability, unlocking broader applications in autonomous driving. Webpage: flymin.github.io/magicdrive-v2/
@@ 1 Introduction
Figure 2: Different from spatial latent [11, 39, 38, 26, 18], spatial-temporal latent (ours) requires spatial-temporal condition injection for geometry controls (we omit the text condition here).
In this paper, we present MagicDrive-V2, a novel approach based on the DiT architecture, designed to tackle the challenges of high-resolution and long street-view video synthesis with precise control. Building upon DiT [29, 46] integrated with a 3D VAE for geometric control, we introduce an MVDiT block for multi-view video generation, and a novel spatial-temporal conditional encoding for latents derived from CogVAE [42], enabling precise management of geometric controls. Additionally, we enhance contextual description supports by efficiently generating captions for each video. To train the model efficiently, we employ a progressive training strategy, transitioning from short to long videos, which helps the model capture intricate details and generalize to complex scenarios. Furthermore, we utilize videos of diverse resolutions and durations to enhance the model’s generalization capability, enabling it to synthesize longer videos than those used in training.
@@
  • We design an efficient framework, MagicDrive-V2, leveraging progressive training to achieve high-quality, high-resolution, and long video generation (up to $848\times 1600\times 6$, 241 frames), significantly superior to prior works.
@@ 2 Related Work
Front View GAIA-1[15] 288$\times$512$\times$1 26
@@
Multi- view MagicDrive [11] 224$\times$400$\times$6 60
@@
Table 1: Comparison of Resolution and Frame Count. We only consider a single inference, since rollout notably degrades quality (see Appendix I). Only support text & image(s) conditions.
@@ 3 Preliminary
$\mathbf{z}_{t}$ $=(1-t)x_{0}+t\epsilon$ (1)
$\mathcal{L}_{CFM}$ $=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}\|v_{\Theta}(\mathbf{z}_{t},t)-\frac{1}{1-t}(\mathbf{z}-\epsilon)\|_{2}^{2}\text{,}$ (2)
where $t\sim\operatorname{lognorm}(0,1)$ is timestep and $v_{\Theta}$ is the model.
@@ 4.2 Architecture for High-Resolution Long Video
As depicted in Figure 3, MagicDrive-V2 utilizes STDiT-3 blocks as per Zheng et al. [46]. The architecture design incorporates two significant modifications. First, to facilitate multi-view generation, the Multi-View DiT (MVDiT) block integrates a cross-view attention layer [11], as shown on the left of Figure 3. Second, given the need for handling multiple control elements, MagicDrive-V2 employs cross-attention [30] for text, boxes, camera views, and trajectories, alongside an addictive branch [43] for maps to inject control signals.
@@ 4.3 Spatial-Temporal Control on 3D VAE
Besides, the issue of temporal misalignment does not exist in ambiguous control conditions (e.g., text), as these conditions are typically employed to describe the entire video. Specifically, the temporal dimension of such conditions is $1$, and they are repeated for $T/f$ times to control the entire video. Consequently, a naive approach to handle such dimension misalignment would be to follow the method of text control, where the temporal dimension is first reduced to $1$ through a temporal encoder and then repeated to $T/f$ for control. However, our experiments reveal that this operation results in trailing issues (as ”Box Reduce” in Figure 10), which we hypothesize are caused by the repeat operation. Therefore, we design a downsampling method that aligns the dimensions while preserving the distinctiveness of temporal information, thereby enabling precise control of geometric information for each frame.
Figure 4: Spatial-Temporal Encoder for Maps (a) and Boxes (b). Our spatial encoding module follows [11], and temporal encoding integrates the downsampling strategy in our 3D VAE [42], resulting in temporally aligned embedding between the control signals and video latents.
@@
All the downsampling ratios are aligned with the adopted 3D VAE [42], i.e., with $8n$ or $18n+$ 1 as inputs, and $2n$ or $12n+$ 1 as outputs, respectively. Section 5.3 holds more ablation studies on different downsampling methods.
@@ 5.1 Experimental Setups
Dataset and Baselines. We employ the nuScenes dataset [4], a prominent dataset for street-view generation [11, 40, 26], to evaluate MagicDrive-V2. We adhere to the official splits, using 700 multi-view videos for training and 150 for validation. Our primary baseline is MagicDrive [11], comprising three models: MagicDrive (16f) for 16-frame video generation; MagicDrive (60f), which extends MagicDrive (16f) to 60 frames, both from Gao et al. [11]; and MagicDrive3D, a 16-frame model from Gao et al. [10]. Moreover, we employed the Waymo Open Dataset [33] to further fine-tune the model, aiming to validate its scalability and to facilitate the generation of videos with diverse environmental styles and different numbers of perspectives. Further details are provided in Appendix D.
@@
Model Setup. In accordance with Section 4.2, we adopt the 3D VAE framework from CogVideoX [42] and train diffusion models for street-view video generation from scratch [^3] For more discussion on this choice, see Section 5.3 and Appendix H.. Initially, the model comprises only spatial blocks for base and control blocks, focusing on image generation. Temporal blocks are incorporated in the second stage, forming the full architecture of MagicDrive-V2. Further training details can be found in Appendix E.
@@ 5.2 Results and Analysis
Generation Quality. MagicDrive-V2 excels in both video and image generation tasks. In video tasks, it significantly reduces FVD compared to MagicDrive (Table 2), due to the DiT architecture enhancing inter-frame consistency and spatial-temporal condition encoding for precise control of object motion and positioning. Further illustrated in Figure 6, MagicDrive-V2 generates high-resolution videos that not only improve quality but also incorporate more intricate details, closely resembling footage captured by real cameras. This enhancement is achieved through our advanced training on variable lengths and resolutions, allowing for more realistic and detailed outputs. We further conduct human evaluation on consistency in Appendix F.
@@
Figure 8: MagicDrive-V2 generates high-resolution (e.g., 424$\times$800 here) street-view videos for 241 frames (i.e., the full length of nuScenes videos, approximately 20 seconds at 12 FPS) with multiple controls (i.e., road map, object boxes, ego trajectory, and text). Notably, the 241-frame length at 424$\times$800 is unseen during training, demonstrating our method’s generalization capability to video length. We annotate the ego-vehicle trajectory and selected objects to aid localization, with same-color boxes denoting the same object. Due to space constraints, the “rainy” example includes only two frames; additional examples can be found in Appendix J and on our project page.
@@ 5.3 Ablation Studies
VAE Comparison for Street Views. Before training the diffusion models, we evaluate the performance of open-source 3D VAEs (i.e., CogVAE [42] and Open-Sora [46]) on street views, in comparison with the 2D SD VAE [30]. As illustrated in Figure 7, CogVAE consistently outperforms its counterparts on reconstruction ability. Besides, as presented in Appendix G, CogVAE demonstrates minimal performance degradation over longer video sequences, making it particularly well-suited for long video generation tasks. Additionally, we found that all VAEs exhibit improved reconstruction capabilities with increasing resolution. This insight is beneficial for advancing our model’s ability to generate high-quality images and videos by focusing on higher-resolution outputs. For more detailed quantitative results and discussion, see Appendix G.
@@ 6 Applications
Fast Generalization on Other Datasets. To evaluate the generalization of MagicDrive-V2, we fine-tuned the Stage 3 model on the Waymo dataset [33], enabling rapid generation of 3-view videos in $1$ day ($1$k+ steps) with strong controllability (examples shown in Figure 5 and more videos can be found on our project page). Moreover, mixed training on Waymo and nuScenes datasets further enhanced the model’s ability to generate videos with varying perspectives, improving overall quality. The final model achieved an FVD of 105.17 on Waymo and 74.30 on nuScenes, surpassing results in Table 2.
@@ 7 Conclusion and Discussion
Note: Our model is capable of generating videos at a resolution of 848$\times$1600 for 241 frames, which is the highest resolution and frame count in the nuScenes dataset. However, the inference cost is currently substantial. Therefore, the primary numerical results in our paper do not utilize this maximum setting. We have included some generated results in Appendix J and on our project website for reference. Future work may focus on further reducing the inference cost.
@@ C Efficiency of Progressive Bootstrap Training
D More Experimental Details
@@
E More Training Details
@@
Inference. By default, images/videos are sampled using Rectified Flow [9] with 30 steps and the classifier-free-guidance (CFG) scale at $2.0$. To support CFG, we randomly drop different conditions at a rate of 15%, including embeddings for text, camera, ego trajectory, and boxes. We follow Gao et al. [11] to use $\{\mathbf{0}\}$ as the null condition for maps in CFG inference. When inferring high-resolution long videos, we also use sequence parallel (Appendix A) to fit in the limited memory of a single GPU.
F Human Evaluation for Multi-frame & Multi-view Consistency
@@
G More Comparison among VAEs
To quantitatively compare the performance of the VAE, we randomly selected two 6-view videos from the nuScenes dataset and used the PSNR metric to evaluate the VAE’s reconstruction ability. Table III presents the results, averaged across the six views. From these results, we observe that CogVAE [42] demonstrates the best reconstruction ability, even surpassing the 2D VAE [30]. Comparing the results from different settings, we find that the current 3D VAEs exhibit good generalization ability for long videos, primarily due to the window-based downsampling techniques [42, 46]. Additionally, we observe that high-resolution content retains a relatively high PSNR after VAE reconstruction, indicating that the current VAEs are more favorable for high-resolution data. This observation also supports our motivation for high-resolution generation.
@@ G More Comparison among VAEs
Table III: VAE Comparison for Street Views. CogVAE [42] and Open-Sora [46] (1.2) are 3D VAEs; SD VAE [30] is 2D VAE, which is also widely adopted by previous street view generation (e.g., [11]). Results are PSNRs calculated through videos from the nuScenes validation set. MagicDrive-V2 adopts CogVAE.
H Reason for Using CogVAE without the Pre-trained Diffusion Model
@@
We conducted a comparative analysis of VAEs, as detailed in Section 5.3 and Appendix G, and found CogVAE [42] to perform well. Given that the VAE determines the upper limit of generation quality, we opted to use CogVAE for video encoding. Notably, CogVideoX [42] employs a novel DiT structure, where each layer’s latent space integrates both video and text condition information. This approach may complicate the design of geometry-related conditions. Furthermore, CogVideoX was not trained in a driving scenario. To eliminate these potential confounding factors, we decided to train the diffusion model from scratch using CogVAE. This strategy allows us to move beyond the constraints of pre-trained models, enabling more flexible modifications to the model architecture to achieve multi-view consistency and spatiotemporal encoding of geometry conditions.
@@
I Single Inference v.s. Rollout Inference
@@
Figure III: Comparison between Rollout for Long Videos (Vista [12]) and Single Inference (our MagicDrive-V2). Although rollout can handle long videos, the quality is significantly degraded. In contrast, our extrapolation maintains high quality in long video generation.
@@
J More Visualization
As said in the “Note”, MagicDrive-V2 is capable of generating $6\times 848\times 1600\times 241$ videos (20s at 12 fps). We include more generated samples in Figure IV-V. Please see the videos on our project page.
Figure IV: We show some frames from the generated $6\times 848\times 1600\times 241$ videos with the same scene configuration (i.e., boxes, maps, cameras, and ego trajectory) but under different weather conditions. Conditions are from the nuScenes validation set.
Figure V: We show some frames from the generated $6\times 848\times 1600\times 241$ videos with the same scene configuration (i.e., boxes, maps, cameras, and ego trajectory) but under different time-of-day conditions. Conditions are from the nuScenes validation set.

### v4

@@ Abstract
The rapid advancement of diffusion models has greatly improved video synthesis, especially in controllable video generation, which is vital for applications like autonomous driving. Although DiT with 3D VAE has become a standard framework for video generation, it introduces challenges in controllable driving video generation, especially for frame-wise geometric control, rendering existing methods ineffective. To address these issues, we propose MagicDrive-V2, a novel approach that integrates the MVDiT block and spatial-temporal conditional encoding to enable multi-view video generation and precise geometric control. Additionally, we introduce an efficient method for obtaining contextual descriptions for videos to support diverse textual control, along with a progressive training strategy using mixed video data to enhance training efficiency and generalizability. Consequently, MagicDrive-V2 enables multi-view driving video synthesis with $\times3$ resolution and 4×4$\times$4 × frame count (compared to current SOTA), rich contextual control, and geometric controls. Extensive experiments demonstrate MagicDrive-V2’s ability, unlocking broader applications in autonomous driving. Project page: flymin.github.io/magicdrive-v2/
@@ 1 Introduction
Figure 2: Different from spatial latent [12, 40, 39, 27, 19], spatial-temporal latent requires spatial-temporal condition injection (ours) for frame-wise geometry controls.
In this paper, we present MagicDrive-V2, a novel approach designed to tackle the challenges of high-resolution and long street-view video synthesis with precise control. Building upon DiT [29, 47] integrated with 3D VAE [43], we propose an MVDiT block for multi-view video generation. Additionally, we introduce a novel spatial-temporal conditional encoding for frame-wise control of spatial-temporal latents, enabling precise geometric management. Additionally, we enhance contextual description supports by efficiently generating captions for each video. To train the model efficiently, we employ a progressive training strategy, transitioning from short to long videos, which helps the model capture intricate details and generalize to complex scenarios. Furthermore, we utilize videos of diverse resolutions and durations to enhance the model’s generalization capability, enabling it to synthesize longer videos than those used in training.
@@
  • We design an efficient framework, MagicDrive-V2, leveraging progressive training to achieve high-quality, high-resolution, and long video generation (up to 848×1600×6848$\times$ 1600$\times$ 6848 × 1600 × 6, 241 frames), significantly superior to prior works.
@@ 2 Related Work
Front
View
GAIA-1[16] 288$\times$512$\times$1 26
@@
Multi-
view
MagicDrive [12] 224$\times$400$\times$6 60
@@
Table 1: Comparison of Resolution and Frame Count. We only consider a single inference, since rollout notably degrades quality (see Appendix K). Only support text & image(s) conditions.
@@ 3 Preliminary
$\mathbf{z}_{t}$ $=t\mathbf{z}_{1}+(1-t)\epsilon$ (1)
$\mathcal{L}_{CFM}$ $=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}\|v_{\Theta}(\mathbf{z}_{t},t)-(\mathbf{z}_{1}-\epsilon)\|_{2}^{2}\text{,}$ (2)
where $\mathbf{z}_{1}$ is the latent of video from the VAE, $t\sim\operatorname{lognorm}(0,1)$ is timestep and $v_{\Theta}$ is the model.
@@ 4.2 Architecture for High-Resolution Long Video
As depicted in Figure 3, MagicDrive-V2 utilizes STDiT-3 blocks as per Zheng et al. [47]. The architecture design incorporates two significant modifications. First, we propose a Multi-View DiT (MVDiT) block, integrating a cross-view attention layer [12] to facilitate multi-view generation, as shown on the left of Figure 3. Second, to handle multiple geometric controls, MagicDrive-V2 utilizes cross-attention [31] for processing text, boxes, camera views, and trajectories, while incorporating an additive branch [44] for maps to inject road structure controls.
@@ 4.3 Spatial-Temporal Control on 3D VAE
Besides, the issue of temporal misalignment does not exist in ambiguous control conditions (e.g., text), as these conditions are typically employed to describe the entire video. Specifically, the temporal dimension of such conditions is 111, and they are repeated for $T/f$ times to control the entire video. Consequently, a naive approach to handle such dimension misalignment would be to follow the method of text control, where the temporal dimension is first reduced to 111 through a temporal encoder and then repeated to $T/f$ for control. However, our experiments reveal that this operation results in trailing issues (as ”Box Reduce” in Figure 10), which we hypothesize are caused by the repeat operation. Therefore, we design a downsampling method that aligns the dimensions while preserving the distinctiveness of temporal information, thereby enabling precise control of geometric information for each frame.
Figure 4: Spatial-Temporal Encoder for Maps (a) and Boxes (b). Our spatial encoding module follows [12], and temporal encoding integrates the downsampling strategy used in our 3D VAE [43], resulting in temporally aligned embedding between the control signals and video latents.
@@
[…]
@@ 5.1 Experimental Setups
Dataset and Baselines. We employ the nuScenes dataset [4], a prominent dataset for street-view generation [12, 41, 27], to evaluate MagicDrive-V2. We adhere to the official splits, using 700 multi-view videos for training and 150 for validation. Our primary baseline is MagicDrive [12], comprising three models: MagicDrive (16f) for 16-frame video generation; MagicDrive (60f), which extends MagicDrive (16f) to 60 frames, both from Gao et al. [12]; and MagicDrive3D, a 16-frame model from Gao et al. [11]. Moreover, we employed the Waymo Open Dataset [34] to further fine-tune the model, aiming to validate its scalability and to facilitate the generation of videos with diverse environmental styles and different numbers of perspectives. Further details are provided in Appendix E.
@@
Model Setup. In accordance with Section 4.2, we adopt the 3D VAE framework from CogVideoX [43] and train diffusion models for street-view video generation from scratch [^3] For more discussion on this choice, see Section 5.3 and Appendix J.. Initially, the model comprises only spatial blocks for base and control blocks, focusing on image generation. Temporal blocks are incorporated in the second stage, forming the full architecture of MagicDrive-V2. Further training details can be found in Appendix F.
@@ 5.2 Results and Analysis
Generation Quality. MagicDrive-V2 excels in both video and image generation tasks. In video tasks, it significantly reduces FVD compared to MagicDrive (Table 2), due to the DiT architecture enhancing inter-frame consistency and spatial-temporal condition encoding for precise control of object motion and positioning. Further illustrated in Figure 6, MagicDrive-V2 generates high-resolution videos that not only improve quality but also incorporate more intricate details, closely resembling footage captured by real cameras. This enhancement is achieved through our advanced training on variable lengths and resolutions, allowing for more realistic and detailed outputs. We further conduct human evaluation on consistency in Appendix G.
@@
Figure 8: MagicDrive-V2 generates high-resolution (e.g., 424$\times$800 here) street-view videos for 241 frames (i.e., the full length of nuScenes videos, approximately 20 seconds at 12 FPS) with multiple controls (i.e., road map, object boxes, ego trajectory, and text). Notably, the 241-frame length at 424$\times$800 is unseen during training, demonstrating our method’s generalization capability to video length. We annotate the ego-vehicle trajectory and selected objects to aid localization, with same-color boxes denoting the same object. Due to space constraints, the “rainy” example includes only two frames; additional examples can be found in Appendix L and on our project page.
@@ 5.3 Ablation Studies
VAE Comparison for Street Views. Before training the diffusion models, we evaluate the performance of open-source 3D VAEs (i.e., CogVAE [43] and Open-Sora [47]) on street views, in comparison with the 2D SD VAE [31]. As illustrated in Figure 7, CogVAE consistently outperforms its counterparts on reconstruction ability. Besides, as presented in Appendix I, CogVAE demonstrates minimal performance degradation over longer video sequences, making it particularly well-suited for long video generation tasks. Additionally, we found that all VAEs exhibit improved reconstruction capabilities with increasing resolution. This insight is beneficial for advancing our model’s ability to generate high-quality images and videos by focusing on higher-resolution outputs. For more detailed quantitative results and discussion, see Appendix I.
@@ 6 Applications
Fast Generalization on Other Datasets. To evaluate the generalization of MagicDrive-V2, we fine-tuned the Stage 3 model on the Waymo dataset [34], enabling rapid generation of 3-view videos in 111 day (111k+ steps) with strong controllability (examples shown in Figure 5 and more videos can be found on our project page). Moreover, mixed training on Waymo and nuScenes datasets further enhanced the model’s ability to generate videos with varying perspectives, improving overall quality. The final model achieved an FVD of 105.17 on Waymo and 74.30 on nuScenes, surpassing results in Table 2.
@@ 7 Conclusion and Discussion
Acknowledgment. This work was supported in part by the Innovation and Technology Fund under Grant No. GHP/141/22SZ, in part by the CUHK Strategic Seed Funding for Collaborative Research Scheme under Grant No. 3136023, and in part by the CUHK Research Matching Scheme under Grant No. 7106937, 8601130, and 8601440. We gratefully acknowledge the support of CANN (Compute Architecture for Neural Networks) and the Ascend AI Processor used for this research. We sincerely thank the following individuals for their assistance in conducting human evaluation of generation quality: Dr. Min Li, Dr. Yijun Yang, Mr. Yingji Zhong, Mr. Zhili Liu, Ms. Xiaomeng Wang, Mr. Pengxiang Li, and others.
@@
Note: Our model is capable of generating videos at a resolution of 848$\times$1600 for 241 frames, which is the highest resolution and frame count in the nuScenes dataset. However, the inference cost is currently substantial. Therefore, the primary numerical results in our paper do not utilize this maximum setting. We have included some generated results in Appendix L and on our project website for reference. Future work may focus on further reducing the inference cost.
@@ C Efficiency of Progressive Bootstrap Training
D Video Generation Speed
Table III shows the breakdown of computation and inference costs, together with comparisons with others. By adopting the sequence parallel, our inference speed is on par with the performance of NVIDIA’s Cosmos-transfer1 [30]. We open-source our implementation and welcome future efforts on optimization.
Method resolution
views
frames
Diff. Steps
(sec/it)
Latent Dec.
Total
Device
MagicDrive-V2
848x1600 3 193 18.03 82.83 11.68 H20
248.24 (1 GPU)
848$\times$1600 6 241 53.74 103.36 28.92
848$\times$1600 6 121 28.18 51.94 8.27
Cosmos-transfer1 [30]
704$\times$1280 6 121 20.88 54 19.92
A100-SXM4
704$\times$1280 1 121 3.48 9 3.32
Table III: Inference speed of MagicDrive-V2. Unless otherwise specified, we use 8 GPUs for testing. As compared with Cosmos-transfer1, our implementation of MagicDrive-V2 offers reasonable inference speeds. We also implemented a parallel decoding strategy for latent decoding, offering 3$\times$ speedup over 1 GPU. Total time contains overhead from CPU offloading. There is no such a model; this row is estimated by 6×6$\times$6 × the single-view time.
E More Experimental Details
@@
F More Training Details
@@
Inference. By default, images/videos are sampled using Rectified Flow [10] with 30 steps and the classifier-free-guidance (CFG) scale at 2.02.02.0. To support CFG, we randomly drop different conditions at a rate of 15%, including embeddings for text, camera, ego trajectory, and boxes. We follow Gao et al. [12] to use $\{\mathbf{0}\}$ as the null condition for maps in CFG inference. When inferring high-resolution long videos, we also use sequence parallel (Appendix A) to fit in the limited memory of a single GPU.
G Human Evaluation for Multi-frame & Multi-view Consistency
@@
Besides, Figure III provides an ablation comparison between with and without the MVDiT block, showing that it is crucial to include such a block for multi-view consistency.
Figure III: Ablation on the MVDiT block. MVDiT block is the key to enabling multi-view consistency in driving video generation.
H Human Evaluation on Text Control
[…]
Figure IV: Confusion matrix to show the text control ability in MagicDrive-V2. MagicDrive-V2 supports the control of 6+ weather conditions.
@@ H Human Evaluation on Text Control
Table IV: VAE Comparison for Street Views. CogVAE [43] and Open-Sora [47] (1.2) are 3D VAEs; SD VAE [31] is 2D VAE, which is also widely adopted by previous street view generation (e.g., [12]). Results are PSNRs calculated through videos from the nuScenes validation set. MagicDrive-V2 adopts CogVAE.
I More Comparison among VAEs
To quantitatively compare the performance of the VAE, we randomly selected two 6-view videos from the nuScenes dataset and used the PSNR metric to evaluate the VAE’s reconstruction ability. Table IV presents the results, averaged across the six views. From these results, we observe that CogVAE [43] demonstrates the best reconstruction ability, even surpassing the 2D VAE [31]. Comparing the results from different settings, we find that the current 3D VAEs exhibit good generalization ability for long videos, primarily due to the window-based downsampling techniques [43, 47]. Additionally, we observe that high-resolution content retains a relatively high PSNR after VAE reconstruction, indicating that the current VAEs are more favorable for high-resolution data. This observation also supports our motivation for high-resolution generation.
J Reason for Using CogVAE without the Pre-trained Diffusion Model
@@ I More Comparison among VAEs
We conducted a comparative analysis of VAEs, as detailed in Section 5.3 and Appendix I, and found CogVAE [43] to perform well. Given that the VAE determines the upper limit of generation quality, we opted to use CogVAE for video encoding. Notably, CogVideoX [43] employs a novel DiT structure, where each layer’s latent space integrates both video and text condition information. This approach may complicate the design of geometry-related conditions. Furthermore, CogVideoX was not trained in a driving scenario. To eliminate these potential confounding factors, we decided to train the diffusion model from scratch using CogVAE. This strategy allows us to move beyond the constraints of pre-trained models, enabling more flexible modifications to the model architecture to achieve multi-view consistency and spatiotemporal encoding of geometry conditions.
@@
K Single Inference v.s. Rollout Inference
@@
Figure V: Comparison between Rollout for Long Videos (Vista [13]) and Single Inference (our MagicDrive-V2). Although rollout can handle long videos, the quality is significantly degraded. In contrast, our extrapolation maintains high quality in long video generation.
@@
L More Visualization
As said in the “Note”, MagicDrive-V2 is capable of generating 6×848×1600×2416$\times$ 848$\times$ 1600$\times$ 2416 × 848 × 1600 × 241 videos (20s at 12 fps). We include more generated samples in Figure VI-VII. Please see the videos on our project page.
Figure VI: We show some frames from the generated 6×848×1600×2416$\times$ 848$\times$ 1600$\times$ 2416 × 848 × 1600 × 241 videos with the same scene configuration (i.e., boxes, maps, cameras, and ego trajectory) but under different weather conditions. Conditions are from the nuScenes validation set.
Figure VII: We show some frames from the generated 6×848×1600×2416$\times$ 848$\times$ 1600$\times$ 2416 × 848 × 1600 × 241 videos with the same scene configuration (i.e., boxes, maps, cameras, and ego trajectory) but under different time-of-day conditions. Conditions are from the nuScenes validation set.